OpenStudy (anonymous):

two lines in a coordinate plane have no points of intersection. Which pair could be the equations of the lines? a. 3x-2y=6 6x-4y=12 b.7x+y=10 3x-2y=-3 c.(3/4)x+y=5 3x+4y=2 d.(1/3)x+y=1 3x+4y=8

OpenStudy (aripotta):

if they have no points of intersection, that would make them parallel lines

OpenStudy (aripotta):

and what do we know about parallel lines and slopes...

OpenStudy (aripotta):

parallel lines have the same slope. so if we put the answer choices into slope-intercept form, we can see which choice has the same slope for each line

OpenStudy (anonymous):

oh see it now

OpenStudy (anonymous):

i meant oh ok i see it now

OpenStudy (aripotta):

so which do you think is the answer?

OpenStudy (anonymous):

give me one sec.....

OpenStudy (anonymous):

well i know the slope equation is rise over run (rise/run)

OpenStudy (aripotta):

we don't need to use that. we just need to change the answer choices from standard form to slope-intercept form

OpenStudy (anonymous):

ok now i reall see it

OpenStudy (aripotta):

lol

OpenStudy (anonymous):

a. y=3x-2y y=6x-4y this is what you are talking about. right?

OpenStudy (anonymous):

this is just an example

OpenStudy (aripotta):

slope-intercept form is y = mx + b. so we're basically solving for y in terms of x. y would be on the left by itself

OpenStudy (anonymous):

if i rewrite the answer choses it would look like y=3x-2 2=2y. 2 is the y- intercept

OpenStudy (aripotta):

no...we're using a as an example, right? lemme show you

OpenStudy (aripotta):